OpenStudy (anonymous):

3 + (-11) finnd the sum

OpenStudy (curry):

-8

OpenStudy (angela210793):

=3-11=-8

OpenStudy (anonymous):

+*-=- 3-11=-8

OpenStudy (lalaly):

its like 11-3 just add the - sign so its -8

OpenStudy (anonymous):

The negative sign is greater than the positive sign, so you know the answer will be a negative number, so the answer is therefore -8

OpenStudy (anonymous):

D; thats not an option

OpenStudy (anonymous):

my bad the 3 was meant to be a 4

OpenStudy (lalaly):

so its -7

OpenStudy (anonymous):

okay thanks